OpenStudy (kirbykirby):

Solve for p1: \[\frac{x_1}{p_1}-\frac{100-\sum_{i=1}^{5}x_i}{1-\sum_{i=1}^{5}p_i}=0\]

OpenStudy (kirbykirby):

We also know the following: \[\sum_{i=1}^{6}p_i=1\implies p_6=1-\sum_{i=1}^{5}p_i\]and \[\sum_{i=1}^{6}x_i=100 \implies x_6=100-\sum_{i=1}^{5}x_i\] The answer given is \[p_1=\frac{x_1}{100}\] but the best I got was \[p_1=\frac{x_1(1-p_2-p_3-p_4-p_5)}{x_6+x_1}\]

OpenStudy (anonymous):

@kirbykirby Sorry friend, it's not mine, but I need more explanation for your first line of logic

OpenStudy (kirbykirby):

Do you mean Solve for p1: \[\sum_{i=1}^{6}p_i=1\implies p_6=1-\sum_{i=1}^{5}p_i\]?

OpenStudy (anonymous):

yeap

OpenStudy (kirbykirby):

\[\sum_{i=1}^{6}p_i=1\] same as \[p_1+p_2+p_3+p_4+p_5+p_6=1\] \[\implies p_6=1-p_1-p_2-p_3-p_4-p_5\] \[\implies p_6=1-\sum_{i_1}^{5}p_i=1-(p_1+p_2+p_3+p_4+p_5)=1-p_1-p_2-p_3-p_4-p_5\]

OpenStudy (anonymous):

Sorry, I try to follow, but it's tooooooooo far above my head.

OpenStudy (anonymous):

Hello :) how about using this rule: if\[\frac{a}{b}=\frac{c}{d}=...\]then\[\frac{a}{b}=\frac{a+c+...}{b+d+...}\]

OpenStudy (kirbykirby):

I tried doing something like that at some point, but I didn't really get anywhere o_O

OpenStudy (anonymous):

\[\frac{x_1}{p_1}=\frac{100-\sum_{i=1}^{5}x_i}{1-\sum_{i=1}^{5}p_i}=\frac{x_6}{p_6}\]

OpenStudy (anonymous):

right ? :)

OpenStudy (kirbykirby):

I tried that too, but I got \[p_1=\frac{x_1 p_6}{x_6} \implies p_1=\frac{x_1(1-p_1-p_2-p_3-p_4-p_5)}{x_6}\] .... after a few steps gave me what I wrote in my first reply :(

OpenStudy (kirbykirby):

I can't even get Wolfram to solve it because the expression is too complicated for wolfram =(

OpenStudy (kirbykirby):

I tried solving for p2...p5 and plugging those values in but it's a GIANT mess ;(

OpenStudy (amistre64):

\[\frac{x_1}{p_1}-\frac{100-\sum_{i=1}^{5}x_i}{1-\sum_{i=1}^{5}p_i}=0\] \[\frac{x_1}{p_1}=\frac{100-\sum_{i=1}^{5}x_i}{1-\sum_{i=1}^{5}p_i}\] \[\frac{x_1(1-\sum_{i=1}^{5}p_i)}{100-\sum_{i=1}^{5}x_i}=p_1\]

OpenStudy (amistre64):

not to sure where this leads .... \[x_1(1-\sum_{i=1}^{5}p_i)=p_1(100-\sum_{i=1}^{5}x_i)\] \[x_1-\sum_{i=1}^{5}x_1p_i=100p_1-\sum_{i=1}^{5}p_1x_i\] \[\sum_{i=1}^{5}-x_1p_i+\sum_{i=1}^{5}p_1x_i=100p_1-x_1\] \[\sum_{i=1}^{5}(p_1x_i-x_1p_i)=100p_1-x_1\] \[\frac{x_1+\sum_{i=1}^{5}(p_1x_i-x_1p_i)}{100}=p_1\]

OpenStudy (anonymous):

i tried to get something like\[\frac{x_1}{p_1}=\frac{x_2}{p_2}=...=\frac{x_6}{p_6}=\frac{x_1+x_2+...+x_6}{p_1+p_2+...+p_6}=100\]but it seems impossible :(

OpenStudy (amistre64):

without knowing any detail about the sequence of x and p, its hard to try to get anything specific

OpenStudy (amistre64):

100+0+0+0+0+0 = 100

OpenStudy (anonymous):

yes, there must be some other information...maybe !

OpenStudy (kirbykirby):

Sorry my computer busted out a bit lol. Thanks for your help guys nevertheless!! Unfortunately there is no other information in the problem :S